OpenStudy (anonymous):

Dx integral of (t^2 + 1)^10 dt from 0 to x =?

OpenStudy (anonymous):

\[ D_x\int_0^xf(t)\;dt \]First use FTC, where \(F'(t)=f(t)\): \[ D_x\left(F(x)-F(0)\right)=D_xF(x)-D_xF(0) \]Remember that \(F(0)\) is a constant.\[ D_xF(x)=F'(x) \]Now if you let \(x=t\) you get\[ F'(x) = f(x) \]

OpenStudy (anonymous):

Thus: \[ D_x\int _a^xf(t)\;dt = f(x) \]Assuming that \(a\) is some constant with respect to \(x\).

OpenStudy (anonymous):

ok..what's FTC? first theorem of calculus?

OpenStudy (anonymous):

Fundamental theorem of calculus. Given \(F'(x) = f(x)\) then: \[ \int_a^bf(x)\;dx = F(b)-F(a) \]

OpenStudy (anonymous):

That is how I did: \[ D_x\color{red}{\int_0^xf(t)\;dt} =D_x\color{red}{( F(x)-F(0))} \]

OpenStudy (anonymous):

and so the answer would be (x^2 + 1)^10

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

By the way, \[ D_x\int_a^{g(x)}f(t)\;dt=D_x[F(g(x))-F(a)]=D_xF(g(x)) = F'(g(x))g'(x) \]

OpenStudy (anonymous):

wait how does the Dx F(a) cancel out...

OpenStudy (anonymous):

Because \(F(a)\) is a constant with respect to \(x\). Derivatives of constants are \(0\).

OpenStudy (anonymous):

i though the integral of f(x) from a to b is F(b)-F(a)

OpenStudy (anonymous):

Yeah, in this case \(b=x\) and \(a=0\).

OpenStudy (anonymous):

but F(0) is 1...

OpenStudy (anonymous):

right? cuz the derivative cancels out the integral and when u plug in 0 as x u get 1

OpenStudy (anonymous):

i mean as t